# Universality of Layer-Level Entropy-Weighted Quantization Beyond Model Architecture and Size

### v1

@@ 5 Benchmarking with MMLU
To evaluate the effectiveness of the EWQ method, benchmarking is performed using the MMLU dataset, which provides accuracy and perplexity metrics for evaluating models on a wide range of tasks. The official Hugging Face dataset cais/mmlu is used for this evaluation, as described in Hendrycks et al. (2020) (Hendrycks et al., 2020). The MMLU dataset consists of question-answer pairs across 57 subjects, including elementary mathematics, U.S. history, computer science, law, and more. Achieving high accuracy on this dataset requires models to demonstrate strong general knowledge and advanced problem-solving skills.
@@ 6.5 Further Analysis of FastEWQ Optimization
Expanding on the results discussed in Table 14, we further analyze the behavior of the FastEWQ optimizer trained on the full dataset. Notably, for 8-bit quantization, FastEWQ closely approximates the selections made by the EWQ method, reinforcing its effectiveness as a lightweight alternative. This alignment is particularly relevant given the role of 8-bit quantization in matrix decomposition for attention computations, as seen in LLM.int8() (Dettmers et al., 2022b).
For models such as Qwen/Qwen2-7B-Instruct, google/gemma-2-9b-it, and microsoft/Phi-3.5-mini-instruct, FastEWQ identifies nearly the same transformer blocks for quantization as EWQ, reflecting prior observations that efficient classifiers can approximate computationally expensive sensitivity analyses (Li et al., 2021). However, for meta-llama/Meta-Llama-3.1-8B-Instruct, architectural variations result in minor deviations in block selection. Despite these differences, the impact on performance remains negligible ($\sim 6\times 10^{-4}$), aligning with findings that minor quantization discrepancies are often absorbed by model redundancy (Xiao et al., 2024).
Beyond accuracy and perplexity, FastEWQ’s $O(1)$ analysis time complexity provides a substantial speedup over iterative EWQ, with at least a 100x efficiency gain. Notably, for google/gemma-2-9b-it, training FastEWQ on 70% of the dataset outperforms the full-dataset variant by selecting three additional blocks for quantization, leading to a 19.77% reduction in model size. This finding supports previous research (Ashkboos et al., 2024), which highlights how subset training mitigates overfitting in quantization controllers.
In comparing 8-bit mixed and fast 8-bit mixed quantization, FastEWQ consistently reproduces EWQ’s behavior across most models, further validating its reliability. The flexibility of this method surpasses that of global quantization, representing only a single point in the broader mixed-precision strategy space (Gong et al., 2019). Consequently, FastEWQ offers an optimal balance between accuracy, perplexity, model size, and computational efficiency, reinforcing its role as the preferred quantization strategy.
@@ 7 Conclusion
We introduced Entropy-Weighted Quantization (EWQ), a novel architecture- and size-agnostic method for post-training quantization of LLMs. By analyzing entropy distributions across transformer blocks, EWQ identifies layers amenable to precision reduction while preserving critical high-entropy components. Our experiments demonstrate that EWQ maintains MMLU accuracy within 0.5% of full-precision models while reducing memory usage by up to 18%—outperforming uniform quantization baselines like GPTQ (Frantar et al., 2022)—and achieves superior perplexity in some cases, a phenomenon attributed to quantization-induced regularization. The method’s effectiveness spans diverse architectures (1.6B to 70B parameters), including LLaMA, Qwen, Phi, and Gemma, proving its universality across model scales and designs.
@@ 7.1 Future Directions
    Architectural Generalization: Adapting EWQ principles to non-transformer architectures (e.g., SSMs (Gu and Dao, 2023), RWKV (Peng et al., 2023)) and multimodal models.
@@
    Precision Frontiers: Exploring sub-4-bit quantization (2-bit, 1.58-bit (Ashkboos et al., 2023)) combined with entropy-aware sparsity, building on sparsity-aware methods like SparseGPT (Frantar and Alistarh, 2023).
@@
    System Integration: Co-designing EWQ with emerging memory technologies (HBM3, CXL) and kernel-level optimizations, inspired by FlashAttention (Dao et al., 2022) and vLLM (Kwon et al., 2023).
Additional opportunities include theoretical investigations into entropy-robustness relationships, extending information-theoretic frameworks like (Shwartz-Ziv and Tishby, 2017), and federated learning applications where FastEWQ’s metadata-driven approach could enable dynamic precision allocation across distributed systems (Kairouz et al., 2021). Finally, integrating EWQ with activation quantization (Yao et al., 2022) and KV cache compression (Liu et al., 2025) could unlock end-to-end efficient inference pipelines for next-generation LLMs.

### v2

@@ 5 Benchmarking with MMLU
To evaluate the effectiveness of the EWQ method, benchmarking is performed using the MMLU dataset, which provides accuracy and perplexity metrics for evaluating models on a wide range of tasks. The official Hugging Face dataset cais/mmlu is used for this evaluation, as described in Hendrycks et al. (2020) [Hendrycks et al., 2020]. The MMLU dataset consists of question-answer pairs across 57 subjects, including elementary mathematics, U.S. history, computer science, law, and more. Achieving high accuracy on this dataset requires models to demonstrate strong general knowledge and advanced problem-solving skills.
@@ 6.5 Further Analysis of FastEWQ Optimization
Expanding on the results discussed in Table 14, we further analyze the behavior of the FastEWQ optimizer trained on the full dataset. Notably, for 8-bit quantization, FastEWQ closely approximates the selections made by the EWQ method, reinforcing its effectiveness as a lightweight alternative. This alignment is particularly relevant given the role of 8-bit quantization in matrix decomposition for attention computations, as seen in LLM.int8() [Dettmers et al., 2022b].
For models such as Qwen/Qwen2-7B-Instruct, google/gemma-2-9b-it, and microsoft/Phi-3.5-mini-instruct, FastEWQ identifies nearly the same transformer blocks for quantization as EWQ, reflecting prior observations that efficient classifiers can approximate computationally expensive sensitivity analyses [Li et al., 2021]. However, for meta-llama/Meta-Llama-3.1-8B-Instruct, architectural variations result in minor deviations in block selection. Despite these differences, the impact on performance remains negligible ($\sim 6\times 10^{-4}$), aligning with findings that minor quantization discrepancies are often absorbed by model redundancy [Xiao et al., 2024].
Beyond accuracy and perplexity, FastEWQ’s $O(1)$ analysis time complexity provides a substantial speedup over iterative EWQ, with at least a 100x efficiency gain. Notably, for google/gemma-2-9b-it, training FastEWQ on 70% of the dataset outperforms the full-dataset variant by selecting three additional blocks for quantization, leading to a 19.77% reduction in model size. This finding supports previous research [Ashkboos et al., 2024], which highlights how subset training mitigates overfitting in quantization controllers.
In comparing 8-bit mixed and fast 8-bit mixed quantization, FastEWQ consistently reproduces EWQ’s behavior across most models, further validating its reliability. The flexibility of this method surpasses that of global quantization, representing only a single point in the broader mixed-precision strategy space [Gong et al., 2019]. Consequently, FastEWQ offers an optimal balance between accuracy, perplexity, model size, and computational efficiency, reinforcing its role as the preferred quantization strategy.
@@ 7 Conclusion
We introduced Entropy-Weighted Quantization (EWQ), a novel architecture- and size-agnostic method for post-training quantization of LLMs. By analyzing entropy distributions across transformer blocks, EWQ identifies layers amenable to precision reduction while preserving critical high-entropy components. Our experiments demonstrate that EWQ maintains MMLU accuracy within 0.5% of full-precision models while reducing memory usage by up to 18%—outperforming uniform quantization baselines like GPTQ [Frantar et al., 2022]—and achieves superior perplexity in some cases, a phenomenon attributed to quantization-induced regularization. The method’s effectiveness spans diverse architectures (1.6B to 70B parameters), including LLaMA, Qwen, Phi, and Gemma, proving its universality across model scales and designs.
@@ 7.1 Future Directions
    Architectural Generalization: Adapting EWQ principles to non-transformer architectures (e.g., SSMs [Gu and Dao, 2023], RWKV [Peng et al., 2023]) and multimodal models.
@@
    Precision Frontiers: Exploring sub-4-bit quantization (2-bit, 1.58-bit [Ashkboos et al., 2023]) combined with entropy-aware sparsity, building on sparsity-aware methods like SparseGPT [Frantar and Alistarh, 2023].
@@
    System Integration: Co-designing EWQ with emerging memory technologies (HBM3, CXL) and kernel-level optimizations, inspired by FlashAttention [Dao et al., 2022] and vLLM [Kwon et al., 2023].
Additional opportunities include theoretical investigations into entropy-robustness relationships, extending information-theoretic frameworks like [Shwartz-Ziv and Tishby, 2017], and federated learning applications where FastEWQ’s metadata-driven approach could enable dynamic precision allocation across distributed systems [Kairouz et al., 2021]. Finally, integrating EWQ with activation quantization [Yao et al., 2022] and KV cache compression [Liu et al., 2025] could unlock end-to-end efficient inference pipelines for next-generation LLMs.